Add the following numbers in 2's complement: -11 +19 using a word length of 6 bits
Answer #1

11 to binary = 001011

One's complement of -11 = 110100

2's complement of -11 = 110100 + 1 = 110101

19 to binary = 010011

2's complement of 19 = 010011

Hence, -11 + 19 = 110101 + 010011

   1 1 0 1 0 1

+ 0 1 0 0 1 1

-----------------------------------------

0 0 1 0 0 0 (Carry = 1)